Use the quadratic formula to solve the following equations. Show all major steps involved and leave solutions in radical form (when applicable).
a: 4x^2 + 5x = 6

Answer by fractalier(6550)

First put it in standard form, the use the quadratic...so we get
4x^2 + 5x = 6
4x^2 + 5x - 6 = 0
and then
x = (-5 ± sqrt(25 + 96)) / 8
x = (-5 ± sqrt(121)) / 8
x = (-5 ± 11) / 8
x = 3/4, -2